# Symmetry-aware spectral bounds for the electronic structure Hamiltonian

### v1

@@ Abstract
Our findings also indicate that although modern tensor compression methods, especially THC, are close to approaching the conventional spectral bound, there remains a lot of room for improvement in reaching the symmetry ($\Delta_{s}/2$) and symmetry-sector ($\Delta_{\mu}/2)$ bounds. Furthermore, while the work by Loaiza et al. [34, 35] has paved the way in proposing symmetry-shifted and block-invariant Hamiltonians as a methodology towards reducing the overall $\ell_{1}$ norms, it remains to be seen whether these methods can saturate either of the two symmetry-aware bounds for large-scale systems. Note that while they do propose a way of building a block-encoded function that is guaranteed to saturate the lower bound [34], the cost of building this function is prohibitive. This leaves significant opportunity for advancements in tensor compression techniques within electronic structure theory, particularly by leveraging the full set of symmetries in various problem settings. While our study has focused on performance limits tied to $\ell_{1}$ norm, exploring other constraints, such as those related to the block-encoding cost, $C_{\mathcal{W}[H]}$, of the Hamiltonian remains critical. Future research should seek to define the optimal trade-off, or Pareto frontier, between these factors for enhancing quantum algorithm efficiency. This exploration highlights multiple promising paths for future enhancements in quantum algorithm capabilities.
Appendix A Exploring the impact of the Hamiltonian Spectrum on
a quantum algorithm’s performance
The development of high-performing quantum algorithms is contingent on the efficient representation of quantum states and operators. Within the context of physics or chemistry-based problems, a Hamiltonian $\hat{H}$ is often used to encode the problem of interest. To provide a unified computational framework that encompasses most well-known fault-tolerant and early fault-tolerant quantum algorithms, the Hamiltonian is typically expressed as a sum of unitary operators $\hat{U}_{k}$ weighted by complex coefficients $c_{k}$,
$\hat{H}=\sum_{k}^{L}c_{k}\hat{U}_{k},\;\;\text{where}\;\;\lambda=\sum_{k}^{L}|c_{k}|,$ (A.1)
is the $\ell_{1}$ norm of these coefficients. Modern quantum algorithms typically utilize one of two Hamiltonian oracle models. The first is the Hamiltonian evolution (HE) model,
$U_{\mathrm{HE}}=e^{i\hat{H}}.$ (A.2)
The second is the Block Encoding (BE) model, where the unitary operator $U_{\mathrm{BE}}$ is defined as:
$U_{\mathrm{BE}}=\begin{pmatrix}\hat{H}/\lambda&*\\ *&*\end{pmatrix}.$ (A.5)
Both of these Hamiltonian oracles play a key role in preparing zero-temperature ground states, preparing finite temperature Gibbs states, estimating Green’s functions and correlation functions, as well as energy and observable estimation.
Throughout this document, we also consider a Hamiltonian $\hat{H}$ that is invariant under the action of a set of symmetries, $\mathbb{S}=\{\hat{S}_{m}\}$, which collectively form a group $G$. The group $G$ includes all individual group elements for each symmetry, as well as any combinations of these symmetries that leave the Hamiltonian invariant. Maschke’s theorem enables a decomposition of the Hilbert space $\mathcal{H}$ into invariant subspaces, known as representations of $G$, which are determined to be irreducible if they lack further invariant subspaces [43, 44]. An invariant Hamiltonian satisfies, $[\hat{g},\hat{H}]=0$, for all $g\in G$. Schur’s lemma may then be used to reveal the block-diagonal structure of $\hat{H}$,
$\hat{H}=\bigoplus_{\boldsymbol{m}}\bigoplus_{n=1}^{\mathrm{deg}\,\boldsymbol{m}}\hat{H}_{n}^{\boldsymbol{m}}\equiv\bigoplus_{\mu}\hat{H}_{\mu},$ (A.6)
In the following subsections, we discuss the importance of the Hamiltonian spectrum, and in particular the spectral range which acts as a key parameter, influencing the performance of a broad range of quantum algorithms that employ either the HE or BE oracles. In the following subsections, we discuss the importance of the Hamiltonian spectrum, and in particular the spectral range which acts as a key parameter, influencing the performance of a broad range of quantum algorithms that employ either the HE or BE oracles.
Hamiltonian Evolution.
When using quantum phase estimation, the Hamiltonian oracle must be normalized so that the eigenvalue spectrum lies between 0 and $2\pi$ to avoid aliasing. This is achieved by using a modified Hamiltonian, $H^{\prime}=c_{1}H+c_{2}\mathbb{1}$, where $c_{1}$ and $c_{2}$ are normalization constants. The choice of these constants depends on the Hamiltonian type and the input state used in the algorithm. The constant $c_{1}$ is crucial as it rescales the eigenvalue spectrum, influencing the required precision for the output phase.
General Hamiltonians. For general bounded Hermitian operators without symmetries, the ideal normalization constants are $c_{1}=\frac{2\pi}{\Delta}$ and $c_{2}=-\frac{2\pi}{\Delta}E_{\mathrm{min}}$, where $\Delta=E_{\mathrm{max}}-E_{\mathrm{min}}$ is the spectral range, and $E_{\mathrm{min}}/E_{\mathrm{max}}$ are the minimum/maximum eigenvalues of the Hamiltonian operator, $\hat{H}$. These normalization constants are suitable for arbitrary input states. Although these eigenvalues are not known in advance, it is possible to use approximate eigenvalues $\tilde{E}_{\mathrm{min}}/\tilde{E}_{\mathrm{max}}$ satisfying the conditions, $E_{\mathrm{min}}\leq\tilde{E}_{\mathrm{min}}$ and $E_{\mathrm{max}}\leq\tilde{E}_{\mathrm{max}}$. As an example, the methods presented in Appendix C and Appendix D may be used to provide numerically tractable estimates of $\tilde{E}_{\mathrm{min}}/\tilde{E}_{\mathrm{max}}$.
Symmetry-Preserving Hamiltonians. If the Hamiltonian preserves a specific symmetry transformation $U_{S}$, as detailed in the main text, different rescaling parameters should be used. Assuming the input state to the QPE algorithm belongs to the symmetry sector $\mu$, the normalization constants should ideally become equal to $c_{1}=\frac{2\pi}{\Delta_{\mu}}$ and $c_{2}=-\frac{2\pi}{\Delta_{\mu}}E^{\mu}_{\mathrm{min}}$, where $\Delta_{\mu}=E^{\mu}_{\mathrm{max}}-E^{\mu}_{\mathrm{min}}$, and $E^{\mu}_{\mathrm{min}}/E^{\mu}_{\mathrm{max}}$ are the minimum/maximum eigenvalues of the Hamiltonian operator $\hat{H}$ in symmetry sector $\mu$. These rescaling constants are suitable for arbitrary input states belonging to symmetry sector $\mu$. As above, the methods presented in Appendix C and Appendix D may be used to provide numerically tractable estimates of $\tilde{E}^{\mu}_{\mathrm{min}}/\tilde{E}^{\mu}_{\mathrm{max}}$ belonging to a given symmetry sector $\mu$.
Block encoding.
Block-encoding oracles have similar applications to HE oracles since these two input models are in fact equivalent if we ignore the overhead that comes from converting between them. Within the context of GSEE, block-encoding is often used as a building block of the qubitization walk operator, defined as $\mathcal{W}=\exp\left(\pm i\sin^{-1}(\hat{H}/\lambda)\right)$. Since the output phase, $\varphi=\pm\sin^{-1}(E/\lambda)$, is limited between $-\pi/2$ and $\pi/2$, aliasing is not a fundamental concern in this context. The primary consideration is ensuring that $\mathcal{W}$ remains unitary for any input state. This implies the Hamiltonian oracle must have an eigenvalue spectrum bounded between -1 and 1.
General Hamiltonians. For general bounded Hermitian operators without any symmetries, the normalization constant $\lambda$ is equal to the $\ell_{1}$ norm in Eqn (A.1). Since the $\ell_{1}$ norm is highly dependent on the specifics of the Hamiltonian LCU expansion (including tensor compression techniques) as well as orbital optimization methods [38, 21, 22], it is possible to provide a lower bound using the spectral norm $\|\hat{H}\|$ (equal to the largest singular value of $\hat{H}$) or half the spectral range, $\frac{\Delta}{2}\leq\|\hat{H}\|$, as highlighted in Theorem 1 [34]. This implies that the rescaling for general Hamiltonians in the block-encoding oracle model is also limited by the spectral range for arbitrary input states.
Symmetry-preserving Hamiltonians. The situation changes when Hamiltonian preserves set of symmetries $\mathbb{S}=\{\hat{S}_{m}\}$ satisfying the conditions $[\hat{S}_{m},\hat{S}_{n}]=[\hat{S}_{m},\hat{H}]=0,\forall\;m,n\in 1,\cdots,|\mathbb{S}|$. As highlighted in [34, 35], it is possible to define two new Hamiltonians, $H_{S}$, and $H_{BI}$, which themselves have different eigenvalue spectra from $\hat{H}$.
Definition A.1 (Symmetry-shifted Hamiltonian).
The symmetry-shifted Hamiltonian is defined as,
$\hat{H}_{\mathrm{S}}=\hat{H}-\hat{f}(\mathbb{S}),$ (A.7)
Definition A.2 (Block-invariant Hamiltonian).
The block-invariant Hamiltonian is defined as,
$\hat{H}_{\mathrm{BI}}=\hat{H}_{\mathrm{S}}-\hat{B}(\hat{f}^{\prime}(\mathbb{S})-f^{\prime}_{\mu}),$ (A.8)
where $\hat{B}$ is a Hermitian operator that obeys the same symmetries as the Hamiltonian, $[\hat{B},\hat{f}^{\prime}(\mathbb{S})]=0$, but it is assumed that it does not generally commute with the Hamiltonian. The operator $\hat{f}^{\prime}(\mathbb{S})$ is another well-behaved function of the Hamiltonian’s symmetries, $[\hat{H},\hat{f}^{\prime}(\mathbb{S})]=0$, obeying the eigenvalue relation, $\hat{f}^{\prime}(\mathbb{S})|\psi_{\mu}\rangle=f^{\prime}_{\mu}|\psi_{\mu}\rangle$.
While the symmetry-shifted Hamiltonian $\hat{H}_{S}$ preserves the eigenvalue spectrum for all possible symmetry-preserving input states apart from symmetry-sector-dependent identity shifts, the block-invariant Hamiltonian $\hat{H}_{\mathrm{BI}}$ preserves the correct eigenvalue spectrum for the symmetry sector $\mu$ only. As a result, it is possible to show that the symmetry sector bound acts as the true lower bound, $\frac{\Delta_{\mu}}{2}\leq\|\hat{H}_{BI}\|$, for arbitrary input states. The details of this inequality are presented explicitly in Appendix B.
Appendix B Proof of symmetry-aware spectral bounds
In the following, we restate a modified version of the Spectral Bound Theorem first given by Loaiza et al. [34]. See 1
Proof.
Let us denote |$\ket{\Psi_{\mathrm{max}}}$ and |$\ket{\Psi_{\mathrm{min}}}$ as the maximum and minimum eigenvectors of $\hat{H}$, such that
[…]
Assuming a decomposition of the Hamiltonian, $\hat{H}=\sum_{n}c_{n}U_{n}$, where each of the unitaries has a spectral norm given by $\|U_{n}\|=1$, we find:
[…]
$\leq\frac{1}{2}\left(|\!\braket{\Psi_{\mathrm{max}}}{\hat{H}}{\Psi_{\mathrm{max}}}|+|\!\braket{\Psi_{\mathrm{min}}}{\hat{H}}{\Psi_{\mathrm{min}}}|\right)$
$\leq\sum_{n}|c_{n}|\|U_{n}\|=\sum_{n}|c_{n}|,$ (B.10)
where the triangle inequality was used in the last line, thereby completing the proof. ∎
The Corollary presented in the main text is derived directly from the observed block-diagonal structure of the Hamiltonian. This structure emerges specifically when the Hamiltonian exhibits a symmetry transformation, denoted as $\hat{U}_{S}$, satisfying the condition $[\hat{H},\hat{U}_{S}]=0$. It can be shown that under such circumstances, the Hamiltonian can be rotated in the block-diagonal form, $\hat{H}=\bigoplus_{\mu}\hat{H}_{\mu},$ where $\hat{H}_{\mu}$ is a sub-block defined with respect to a particular symmetry sector $\mu$. See 2
Proof.
Note that the original spectral bound can be expressed as:
$\frac{\Delta}{2}=\frac{1}{2}\left(\max_{\mu}E_{\mathrm{max}}^{\mu}-\min_{\mu}E^{\mu}_{\mathrm{min}}\right).$ (B.11)
$\Delta$ $=\max_{\mu}E_{\mathrm{max}}^{\mu}-\min_{\mu}E^{\mu}_{\mathrm{min}}$ (B.12)
$=\max_{\mu}E_{\mathrm{max}}^{\mu}-\theta$ (B.13)
$=\max_{\mu}(E_{\mathrm{max}}^{\mu}-\theta)$ (B.14)
$\Delta_{s}\equiv\max_{\mu}(E_{\mathrm{max}}^{\mu}-E_{\mathrm{min}}^{\mu})\leq\Delta.$ (B.15)
While this Corollary proves that a hierarchy of spectral bounds exists, it does not prove that a symmetry-shifted $\hat{H}_{\mathrm{S}}$ or block-invariant $\hat{H}_{\mathrm{BI}}$ Hamiltonian must have corresponding $\ell_{1}$ norms that are restricted by either of these two bounds. The following two theorems prove that these lower bounds apply directly to symmetry-shifted and block-invariant Hamiltonians introduced in refs. [34, 35].
Theorem 3 (Symmetry-shifted 1-norm bound).
For a bounded symmetry-shifted Hermitian operator, $\hat{H}_{\mathrm{S}}$, all of its possible LCU decompositions will have an associated $\ell_{1}$ norm $\lambda_{\mathrm{S}}$ which is lower bounded by the spectral symmetry bound, $\Delta_{s}/2\equiv\tfrac{1}{2}\max_{\mu}(E_{\mathrm{max}}^{\mu}-E_{\mathrm{min}}^{\mu})$,
$\tfrac{1}{2}\Delta_{s}\leq\lambda_{\mathrm{S}}.$ (B.16)
Proof.
$\frac{1}{2}\left(\max_{\mu}(E_{\mathrm{max}}^{\mu}-f_{\mu})-\min_{\mu}(E^{\mu}_{\mathrm{min}}-f_{\mu})\right)\leq\lambda_{\mathrm{S}}.$ (B.17)
$\frac{1}{2}\left(\max_{\mu}(E_{\mathrm{max}}^{\mu}-f_{\mu})-\min_{\mu}(E^{\mu}_{\mathrm{min}}-f_{\mu})\right)$ $=\frac{1}{2}\left(\max_{\mu}(E_{\mathrm{max}}^{\mu}-f_{\mu})+\max_{\mu}(f_{\mu}-E^{\mu}_{\mathrm{min}})\right)$ (B.18)
$\geq\frac{1}{2}\max_{\mu}\left(E_{\mathrm{max}}^{\mu}-f_{\mu}+f_{\mu}-E^{\mu}_{\mathrm{min}}\right).$ (B.19)
Theorem 4 (Block-invariant 1-norm bound).
Consider the block-invariant Hermitian operator, $\hat{H}_{\mathrm{BI}}$, defined in (A.8) with an optimal symmetry-shifted Hamiltonian $\hat{H}_{\mathrm{S}}$. All of its possible LCU decompositions will have an associated $\ell_{1}$ norm $\lambda_{\mathrm{BI}}$ which is lower bounded by the symmetry sector bound, $\Delta_{\mu}/2\equiv\tfrac{1}{2}(E_{\mathrm{max}}^{\mu}-E_{\mathrm{min}}^{\mu})$,
$\tfrac{1}{2}\Delta_{\mu}\leq\lambda_{\mathrm{BI}}.$ (B.20)
Proof.
The first term in (A.8) corresponds to the symmetry-shifted Hamiltonian. If we assume that the symmetry-shifted Hamiltonian has been optimized so that it saturates its bound, $\tfrac{1}{2}\Delta_{s}$, then it is possible to show that $\lambda_{\mathrm{BI}}$ can become smaller than $\tfrac{1}{2}\Delta_{s}$ for a certain set of operators $\hat{B}$ and $\hat{f}^{\prime}(S)$. This illustrates that $\tfrac{1}{2}\Delta_{s}$ is not a proper lower bound. Nevertheless, Corollary 2 shows that $\lambda_{\mathrm{BI}}$ must remain lower bounded by the symmetry sector bound, $\tfrac{1}{2}\Delta_{\mu}$, regardless of the chosen operators $\hat{B}$ and $\hat{F}^{\prime}$. This completes the proof. ∎
Appendix C Symmetry Sector Scaling Analysis
C.1 Linear Combination of Hermitian Operators
@@
$\tfrac{1}{2}\Delta\leq\sum_{i}\tfrac{1}{2}\Delta^{(i)},$ (C.21)
@@
$\tfrac{1}{2}\Delta_{\mu}\leq\sum_{i}\tfrac{1}{2}\Delta^{(i)}_{\mu}.$ (C.22)
@@
C.2 Application: electronic structure Hamiltonian
To elucidate the scaling behavior of the electronic structure Hamiltonian, we use the standard Majorana representation of the Hamiltonian [21, 38],
$\hat{H}=\mathcal{E}_{o}+H_{\mathrm{1body}}+H_{\mathrm{2body}},$ (C.23)
@@ C.2 Application: electronic structure Hamiltonian
$H_{\mathrm{1body}}$ $=\frac{i}{2}\sum_{\begin{subarray}{c}pq\\ \sigma\end{subarray}}\kappa_{pq}\hat{\gamma}_{p\sigma,0}\hat{\gamma}_{q\sigma,1},$ (C.26)
$H_{\mathrm{2body}}$ $=-\frac{1}{8}\sum_{\begin{subarray}{c}pqrs\\ \sigma\tau\end{subarray}}g_{pqrs}\hat{\gamma}_{p\sigma,0}\hat{\gamma}_{q\sigma,1}\hat{\gamma}_{r\tau,0}\hat{\gamma}_{s\tau,1},$ (C.29)
@@
$\mathcal{E}_{o}$ $=\left(E_{\mathrm{nuc}}+\sum_{p}f_{pp}+\tfrac{1}{2}\sum_{pr}g_{pprr}-\tfrac{1}{2}\sum_{pr}g_{prrp}\right),$ (C.30)
$\kappa_{pq}$ $=\left(f_{pq}-\frac{1}{2}\sum_{r}g_{prrq}+\sum_{r}g_{pqrr}\right).$ (C.31)
@@
$\tfrac{1}{2}\Delta\leq\tfrac{1}{2}\Delta_{\mathrm{incoh.}}\equiv\tfrac{1}{2}\Delta^{\mathrm{1body}}+\tfrac{1}{2}\Delta^{\mathrm{2body}},$ (C.32)
@@
$H_{\mathrm{1body}}=\frac{i}{2}\sum_{\begin{subarray}{c}pq\\ \sigma\end{subarray}}\kappa_{pq}\hat{\gamma}_{p\sigma,0}\hat{\gamma}_{q\sigma,1}=\frac{i}{2}\hat{G}^{\dagger}\left(\sum_{k,\sigma}\lambda_{k}\hat{\gamma}^{\dagger}_{k\sigma,0}\hat{\gamma}_{k\sigma,1}\right)\hat{G},$ (C.34)
@@
$\frac{1}{2}\Delta_{\mu}^{\mathrm{1body}}=\frac{1}{2}\left(\sum_{\begin{subarray}{c}k=N_{\mathrm{orb}}-\eta_{\sigma}\\ \sigma\end{subarray}}^{N_{\mathrm{orb}}-1}\!\!\!\!\lambda_{k}-\sum_{\begin{subarray}{c}k=0\\ \sigma\end{subarray}}^{\eta_{\sigma}-1}\lambda_{k}\right).$ (C.39)
@@
$\|H_{\mathrm{1body}}\|_{\mu}=\sum_{\sigma}\text{max}\left[\Bigg{|}\frac{1}{2}\!\!\sum_{k=0}^{N_{\mathrm{orb}}-1}\!\!\!\lambda_{k}-\sum_{\begin{subarray}{c}k=0\end{subarray}}^{\eta_{\sigma}-1}\lambda_{k}\Bigg{|},\Bigg{|}\frac{1}{2}\!\!\sum_{k=0}^{N_{\mathrm{orb}}-1}\!\!\!\lambda_{k}-\!\!\!\!\!\!\sum_{\begin{subarray}{c}k=N_{\mathrm{orb}}-\eta_{\sigma}\end{subarray}}^{N_{\mathrm{orb}}-1}\!\!\!\!\lambda_{k}\Bigg{|}\right],$ (C.40)
which is equal to the largest singular value of $H_{\mathrm{1body}}$ in the symmetry sector $\mu=(\eta_{\alpha},\eta_{\beta}$). The symmetry sector spectral gap (C.39) is a lower bound to the fermionic semi-norm, $\frac{1}{2}\Delta_{\mu}^{\mathrm{1body}}\leq\|H_{\mathrm{1body}}\|_{\mu}$, which is not saturated for the 1-body Hamiltonian expressed in the form above. It is important to note that these results provide a simple and direct way of computing the spectral gaps exactly for 1-body Hamiltonian based on $\{\hat{N},\hat{S}_{z}\}$ symmetries. To elucidate the scaling behavior, we use the following bound
$\|H_{\mathrm{1body}}\|_{\mu}\leq\|\kappa_{pq}\|(\eta_{\alpha}+\eta_{\beta})=\text{max}[|\Lambda|]\eta,$ (C.41)
@@
$g_{pqrs}=\sum_{tkl}\alpha^{t}_{k}\alpha^{t}_{l}\;U^{(t)}_{pk}U^{(t)}_{qk}U^{(t)}_{rl}U^{(t)}_{sl},$ (C.42)
which transforms the two-body Hamiltonian (C.31) into the complete-square form,
$H_{\mathrm{2body}}=\frac{1}{2}\sum_{t}\hat{v}^{2}_{t}$ (C.43)
@@
$\hat{v}_{t}=\hat{G}_{t}^{\dagger}\left(\frac{i}{2}\sum_{\begin{subarray}{c}k\\ \sigma\end{subarray}}\alpha^{t}_{k}\hat{\gamma}_{k\sigma,0}\hat{\gamma}_{k\sigma,1}\right)\hat{G}_{t},$ (C.44)
@@
$\tfrac{1}{2}\Delta_{\mu}^{\mathrm{2body}}\leq\tfrac{1}{4}\sum_{t}\Delta_{\mu}^{(v^{2}_{t})}$ (C.45)
@@
$\Delta_{\mu}^{(v^{2}_{t})}$ $=(E_{\mathrm{max}}^{(\mu,v^{2}_{t})}-E_{\mathrm{min}}^{(\mu,v^{2}_{t})})$ (C.46)
$\leq\|\hat{v}_{t}^{2}\|_{\mu}=\|\hat{v}_{t}\|_{\mu}^{2}.$ (C.47)
@@
$\|\hat{v}_{t}\|_{\mu}=\sum_{\sigma}\text{max}\left[\Bigg{|}\frac{1}{2}\sum_{k=0}^{N_{\mathrm{orb}}-1}\alpha^{t}_{k}-\sum_{\begin{subarray}{c}k=0\end{subarray}}^{\eta_{\sigma}-1}\alpha^{t}_{k}\Bigg{|},\Bigg{|}\frac{1}{2}\sum_{k=0}^{N_{\mathrm{orb}}-1}\alpha^{t}_{k}-\!\!\!\!\!\!\sum_{\begin{subarray}{c}k=N_{\mathrm{orb}}-\eta_{\sigma}\end{subarray}}^{N_{\mathrm{orb}}-1}\!\!\!\!\alpha^{t}_{k}\Bigg{|}\right],$ (C.48)
@@
$\tfrac{1}{2}\Delta_{\mu}^{\mathrm{2body}}\leq\tfrac{1}{4}\sum_{t}\|\hat{v}_{t}\|^{2}_{\mu}\leq\frac{1}{4}\sum_{t}\mathcal{O}\left(g(N_{\mathrm{orb}})\eta^{2}\right),$ (C.49)

### v2

@@ VI Conclusion
Our findings also indicate that although modern tensor compression methods, especially THC, are close to approaching the conventional spectral bound $\Delta/2$, there remains a lot of room for improvement in reaching the symmetry ($\Delta_{s}/2$) and symmetry-sector ($\Delta_{\mu}/2)$ bounds. Loaiza et al. [34, 35] proposed symmetry-shifted and block-invariant Hamiltonians as a methodology for reducing the 1-norms. Recent work by Rocca et al. [rocca2024reducing] exploited such symmetry shifts to reduce the 1-norm towards the symmetry bound $\Delta_{s}/2$ for FeMoco and the P450. However, there is still significant opportunity for advancements in tensor compression techniques, in particular by leveraging the full set of symmetries to reach 1-norms saturating the symmetry-sector bound $\Delta_{\mu}/2$. While our study has focused on performance limits tied to 1-norm, exploring other constraints, such as those related to the block-encoding cost, $C_{\mathcal{W}[H]}$, of the Hamiltonian remains critical. Future research should seek to define the optimal trade-off, or Pareto frontier, between these factors for enhancing quantum algorithm efficiency. This exploration highlights multiple promising paths for future enhancements in quantum algorithm capabilities.
Appendix A Symmetry Sector Scaling Analysis
A.1 Linear Combination of Hermitian Operators
@@
$\tfrac{1}{2}\Delta\leq\sum_{i}\tfrac{1}{2}\Delta^{(i)},$ (A.1)
@@
$\tfrac{1}{2}\Delta_{\mu}\leq\sum_{i}\tfrac{1}{2}\Delta^{(i)}_{\mu}.$ (A.2)
@@
A.2 Application: electronic structure Hamiltonian
To elucidate the scaling behavior of the electronic structure Hamiltonian, we use the standard Majorana representation [21, 38],
$\hat{H}=\mathcal{E}_{o}+H_{\mathrm{1body}}+H_{\mathrm{2body}},$ (A.3)
@@ A.2 Application: electronic structure Hamiltonian
$H_{\mathrm{1body}}$ $=\frac{i}{2}\sum_{\begin{subarray}{c}pq\\ \sigma\end{subarray}}\kappa_{pq}\hat{\gamma}_{p\sigma,0}\hat{\gamma}_{q\sigma,1},$ (A.4)
$H_{\mathrm{2body}}$ $=-\frac{1}{8}\sum_{\begin{subarray}{c}pqrs\\ \sigma\tau\end{subarray}}g_{pqrs}\hat{\gamma}_{p\sigma,0}\hat{\gamma}_{q\sigma,1}\hat{\gamma}_{r\tau,0}\hat{\gamma}_{s\tau,1},$ (A.5)
@@
$\mathcal{E}_{o}$ $=E_{\mathrm{nuc}}+\sum_{p}h_{pp}+\tfrac{1}{2}\sum_{pr}g_{pprr}-\tfrac{1}{2}\sum_{pr}g_{prrp},$ (A.6)
$\kappa_{pq}$ $=h_{pq}-\frac{1}{2}\sum_{r}g_{prrq}+\sum_{r}g_{pqrr}.$ (A.7)
@@
$\tfrac{1}{2}\Delta\leq\tfrac{1}{2}\Delta_{\mathrm{incoh.}}\equiv\tfrac{1}{2}\Delta^{\mathrm{1body}}+\tfrac{1}{2}\Delta^{\mathrm{2body}},$ (A.8)
@@
$H_{\mathrm{1body}}=\frac{i}{2}\sum_{\begin{subarray}{c}pq\\ \sigma\end{subarray}}\kappa_{pq}\hat{\gamma}_{p\sigma,0}\hat{\gamma}_{q\sigma,1}=\frac{i}{2}\hat{G}^{\dagger}\left(\sum_{k,\sigma}\lambda_{k}\hat{\gamma}^{\dagger}_{k\sigma,0}\hat{\gamma}_{k\sigma,1}\right)\hat{G},$ (A.10)
@@
$\frac{1}{2}\Delta_{\mu}^{\mathrm{1body}}=\frac{1}{2}\left(\sum_{\begin{subarray}{c}k=N_{\mathrm{orb}}-\eta_{\sigma}\\ \sigma\end{subarray}}^{N_{\mathrm{orb}}-1}\!\!\!\!\lambda_{k}-\sum_{\begin{subarray}{c}k=0\\ \sigma\end{subarray}}^{\eta_{\sigma}-1}\lambda_{k}\right).$ (A.11)
@@
$\|H_{\mathrm{1body}}\|_{\mu}=\sum_{\sigma}\text{max}\left[\Bigg{|}\frac{1}{2}\!\!\sum_{k=0}^{N_{\mathrm{orb}}-1}\!\!\!\lambda_{k}-\sum_{\begin{subarray}{c}k=0\end{subarray}}^{\eta_{\sigma}-1}\lambda_{k}\Bigg{|},\Bigg{|}\frac{1}{2}\!\!\sum_{k=0}^{N_{\mathrm{orb}}-1}\!\!\!\lambda_{k}-\!\!\!\!\!\!\sum_{\begin{subarray}{c}k=N_{\mathrm{orb}}-\eta_{\sigma}\end{subarray}}^{N_{\mathrm{orb}}-1}\!\!\!\!\lambda_{k}\Bigg{|}\right],$ (A.12)
which is equal to the largest singular value of $H_{\mathrm{1body}}$ in the symmetry sector $\mu=(\eta_{\alpha},\eta_{\beta}$). The symmetry sector spectral gap (A.11) is a lower bound to the fermionic semi-norm, $\frac{1}{2}\Delta_{\mu}^{\mathrm{1body}}\leq\|H_{\mathrm{1body}}\|_{\mu}$, which is not saturated for the 1-body Hamiltonian expressed in the form above. It is important to note that these results provide a simple and direct way of computing the spectral gaps exactly for 1-body Hamiltonian based on $\{\hat{N},\hat{S}_{z}\}$ symmetries. To elucidate the scaling behavior, we use the following bound
$\|H_{\mathrm{1body}}\|_{\mu}\leq\|\kappa_{pq}\|(\eta_{\alpha}+\eta_{\beta})=\text{max}[|\Lambda|]\eta,$ (A.13)
@@
$g_{pqrs}=\sum_{tkl}\alpha^{t}_{k}\alpha^{t}_{l}\;U^{(t)}_{pk}U^{(t)}_{qk}U^{(t)}_{rl}U^{(t)}_{sl},$ (A.14)
which transforms the two-body Hamiltonian (A.7) into the complete-square form,
$H_{\mathrm{2body}}=\frac{1}{2}\sum_{t}\hat{v}^{2}_{t}$ (A.15)
@@
$\hat{v}_{t}=\hat{G}_{t}^{\dagger}\left(\frac{i}{2}\sum_{\begin{subarray}{c}k\\ \sigma\end{subarray}}\alpha^{t}_{k}\hat{\gamma}_{k\sigma,0}\hat{\gamma}_{k\sigma,1}\right)\hat{G}_{t},$ (A.16)
@@
$\tfrac{1}{2}\Delta_{\mu}^{\mathrm{2body}}\leq\tfrac{1}{4}\sum_{t}\Delta_{\mu}^{(v^{2}_{t})}$ (A.17)
@@
$\Delta_{\mu}^{(v^{2}_{t})}$ $=(E_{\mathrm{max}}^{(\mu,v^{2}_{t})}-E_{\mathrm{min}}^{(\mu,v^{2}_{t})})$ (A.18)
$\leq\|\hat{v}_{t}^{2}\|_{\mu}=\|\hat{v}_{t}\|_{\mu}^{2}.$ (A.19)
@@
$\|\hat{v}_{t}\|_{\mu}=\sum_{\sigma}\text{max}\left[\Bigg{|}\frac{1}{2}\sum_{k=0}^{N_{\mathrm{orb}}-1}\alpha^{t}_{k}-\sum_{\begin{subarray}{c}k=0\end{subarray}}^{\eta_{\sigma}-1}\alpha^{t}_{k}\Bigg{|},\Bigg{|}\frac{1}{2}\sum_{k=0}^{N_{\mathrm{orb}}-1}\alpha^{t}_{k}-\!\!\!\!\!\!\sum_{\begin{subarray}{c}k=N_{\mathrm{orb}}-\eta_{\sigma}\end{subarray}}^{N_{\mathrm{orb}}-1}\!\!\!\!\alpha^{t}_{k}\Bigg{|}\right],$ (A.20)
@@
$\tfrac{1}{2}\Delta_{\mu}^{\mathrm{2body}}\leq\tfrac{1}{4}\sum_{t}\|\hat{v}_{t}\|^{2}_{\mu}\leq\frac{1}{4}\sum_{t}\mathcal{O}\left(g(N_{\mathrm{orb}})\eta^{2}\right),$ (A.21)